# Infer-and-widen versus split-and-condition: two tales of selective inference

### v1

@@ Abstract
Recent attention has focused on the development of methods for post-selection inference. However, the connections between these methods, and the extent to which one might be preferred to another, remain unclear. In this paper, we classify existing methods for post-selection inference into one of two frameworks: infer-and-widen or split-and-condition. The infer-and-widen framework produces confidence intervals whose midpoints are biased due to selection, and must be wide enough to account for this bias. By contrast, split-and-condition directly adjusts the intervals’ midpoints to account for selection. We compare the two frameworks in three vignettes: the winner’s curse, maximal contrasts, and inference after the lasso. Our results are striking: in each of these examples, a split-and-condition strategy leads to confidence intervals that are much narrower than the state-of-the-art infer-and-widen proposal, when methods are tuned to yield identical selection events. Furthermore, even an “oracle” infer-and-widen confidence interval — the narrowest possible interval that could be theoretically attained via infer-and-widen — is not necessarily narrower than a feasible split-and-condition method. Taken together, these results point to split-and-condition as the most promising framework for post-selection inference in real-world settings.
@@ 1 Introduction
In classical statistics, the target of inference — for instance, the null hypothesis that we wish to test, or the parameter for which we wish to construct a confidence interval — is fixed. However, in recent years, and motivated by the explosion of data across a variety of fields, interest has focused on the setting where the target of inference is itself a function of the data, and therefore is random. The terms selective inference and post-selection inference are interchangeably used to describe the set of strategies available for conducting inference on a random target. In this paper, we categorize the existing approaches for selective inference into one of two frameworks, which we will refer to as split-and-condition and infer-and-widen. These two frameworks are outlined in Boxes 1 and 1.
The simplest example of the split-and-condition framework (Box 1), sample splitting, involves splitting the data into independent training and test sets. The training set can be used to define the target of inference, and the test set to conduct inference on this target. Because the two sets are independent, we can conduct inference on the test set as though the target of inference were fixed, despite the fact that it is actually a function of the training data: thus, despite the name “split-and-condition”, no conditioning is required. Sample splitting provides a simple and attractive approach to generate independent training and test sets that is applicable in the case of independent and identically distributed (i.i.d.) observations (Cox, 1975). More recent data thinning proposals enable the generation of independent training and test sets in the absence of i.i.d. observations, and even when just a single observation is available (Rasines and Young, 2023; Neufeld et al., 2024; Dharamshi et al., 2024).
The split-and-condition strategy can also be applied in settings when it is either impossible or undesirable to create independent training and test sets. For instance, a researcher may wish to define the target of inference using all of the data. In this case, they can conduct inference conditional on the event that the data led to this particular target of inference, as in the conditional selective inference proposals of Fithian et al. (2017); Lee et al. (2016); Chen and Witten (2023); Gao et al. (2024). In this case, we can think of the data as having been trivially “split” into identical training and test sets that contain the entire data. Alternatively, if a researcher is willing to define the target of inference using only a subset of the data, but no strategy is available to create independent training and test sets, then they might split the data into dependent training and test sets; a target of inference can then be defined on the training set, and inference conditional on the training set can be conducted on the test set, as in the data fission proposal of Leiner et al. (2023). Although conditional selective inference and data fission provide conditional (rather than marginal) inferential guarantees, conditional guarantees imply marginal guarantees via the law of total expectation.
[!htb] Box 1: Split-and-condition. 1. The data is split into (possibly overlapping, or even identical) training and test sets. 2. The target of inference is selected on the training set. 3. Inference is conducted on the test set, conditional on the target of inference selected on the training set.
In Section 2.1, we will elaborate on the split-and-condition framework laid out in Box 1, and will clarify how each of the methods discussed in the preceding paragraphs fits into this framework.
By contrast, the infer-and-widen framework, presented in Box 1, involves (Step 1) defining a target of inference using some or all of the data, and then (Step 2) adjusting the critical values arising from “classical” inference to account for the fact that the target of inference is not fixed. The simplest examples of this strategy are the Bonferroni (Bonferroni, 1936), Holm-Bonferroni (Holm, 1979), and Šidák (Šidák, 1967) procedures, which are suitable when a finite number of targets of inference are considered (and, in the case of Šidák, under additional assumptions). In the special case of inference on contrasts of a linear model, the Scheffé correction is available (Scheffé, 1959); this idea is further refined by the PoSI method (Berk et al., 2013). While conceptually attractive, these approaches tend to be quite conservative, and require consideration of a finite number of targets of inference (Bonferroni, Holm-Bonferroni, Šidák), a linear model setting (Scheffé, PoSI), or are computationally intractable (PoSI).
The algorithmic stability proposal of Zrnic and Jordan (2023) is an elegant recent addition to the infer-and-widen literature. By drawing connections with the field of differential privacy (Dwork and Roth, 2014), the “stability” of the approach for selecting the target of inference in Step 1 of Box 1 is quantified; this, in turn, leads to a stability-specific adjustment of the critical values in Step 2 of Box 1. In a sense, algorithmic stability achieves the lofty goals laid out in the PoSI proposal of Berk et al. (2013) by creating a framework to compute a selection-specific adjustment to the critical values; Zrnic and Jordan (2023) further argue that algorithmic stability has higher power than sample splitting in some settings. We discuss algorithmic stability in greater detail in Section 2.2.
[!htb] Box 2: Infer-and-widen. 1. A target of inference is selected on all of the data. 2. “Classical” inference is conducted on all of the data, with adjustment of the critical values to account for the data-driven selection of the target of inference.
Our paper has three main contributions.
  1. 1.
    We introduce the split-and-condition and infer-and-widen frameworks, which provide a lens through which to view existing proposals for selective inference.
  2. 2.
    We compare these two frameworks within the context of two vignettes that were previously studied in Zrnic and Jordan (2023) in the context of algorithmic stability. In each vignette, we develop a split-and-condition method that uses exactly the same selection event as the algorithmic stability proposal of Zrnic and Jordan (2023). In both vignettes, the former yields subtantially narrower confidence intervals than the latter.
  3. 3.
    Finally, we consider a third vignette involving the lasso (Tibshirani, 1996), for which no infer-and-widen approach is as-of-yet available without extremely over-conservative correction (Bonferroni, 1936) or substantial modification to the underlying algorithm (Zrnic and Jordan, 2023). We show that a simple split-and-condition approach is competitive with an “oracle” infer-and-widen procedure, which is not attainable in practice as it relies on knowledge of the true parameter value.
The bottom line is as follows: when we develop split-and-condition strategies and infer-and-widen strategies that make use of identical selection events, the former outperform the latter. Furthermore, the former often outperform even the “oracle” infer-and-widen procedure.
The rest of this paper is organized as follows. In Section 2, we discuss the split-and-condition and infer-and-widen frameworks in greater detail. In Sections 3, 4, and 5, we compare these two frameworks in two vignettes considered in Zrnic and Jordan (2023), and in an example involving model selection via the lasso. In Section 6, we show analytically that the performance of the infer-and-widen interval is due to bias in the interval midpoint. We close with a discussion in Section 7. Theoretical results are proven in the Appendix. Scripts to fully reproduce all numerical results can be found at
https://github.com/rflperry/two_tales_of_selective_inference.
2 Strategies for selective inference
@@
Consider a random variable $Y$ with domain $\mathcal{Y}$, and a collection of parameters $\theta$ whose elements can be indexed by a (possibly infinite) index set $\mathcal{I}$. For a fixed index $i\in\mathcal{I}$, and for $\alpha\in(0,1)$, we let $\mathrm{CI}_{i}^{\alpha}(Y)$ denote the confidence interval for $\theta_{i}$, derived from the distribution of $Y$, with $1-\alpha$ coverage; that is,
@@
In what follows, we will refer to $\mathrm{CI}_{i}^{\alpha}(Y)$ in (2.1) as a classical confidence interval.
@@
Now, suppose that $i\in\mathcal{I}$ is no longer a fixed index; instead, some selection rule $S:\mathcal{Y}\to\mathcal{I}$ uses the data to select a parameter on which to conduct inference. It has been well-established that (2.1) does not provide a valid $1-\alpha$ confidence interval for $\theta_{\mathcal{S}(Y)}$; that is, $\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{\alpha}(Y)\right)$ may far exceed $\alpha$. The goal of selective inference is to enable valid inference on $\theta_{\mathcal{S}(Y)}$ (Fithian et al., 2017; Gao et al., 2024; Chen and Witten, 2023; Chen and Bien, 2020; Neufeld et al., 2022; Lee et al., 2016; Hyun et al., 2018, 2021; Jewell et al., 2022).
We now will categorize the existing selective inference approaches as either split-and-condition or infer-and-widen.
2.1 The split-and-condition framework
Box 1 outlines the split-and-condition framework. In Step 1, we let $Y^{\mathrm{train}}$ and $Y^{\mathrm{test}}$ denote the training and test sets. In Step 2, the target of inference is $\theta_{\mathcal{S}(Y^{\mathrm{train}})}$. In Step 3, we construct a confidence interval $\mathrm{CI}_{\mathcal{S}(Y^{\mathrm{train}})}^{\alpha}\left(Y^{\mathrm{test}}\mid\mathcal{S}(Y^{\mathrm{train}})\right)$ that satisfies the conditional coverage guarantee
$\mathbb{P}\left(\theta_{\mathcal{S}(Y^{\mathrm{train}})}\not\in\mathrm{CI}_{\mathcal{S}(Y^{\mathrm{train}})}^{\alpha}\left(Y^{\mathrm{test}}\mid\mathcal{S}(Y^{\mathrm{train}})\right)\mid\mathcal{S}(Y^{\mathrm{train}})\right)\leq\alpha.$ (2.2)
However, it is sometimes more convenient to construct a confidence interval
$\mathrm{CI}_{\mathcal{S}(Y^{\mathrm{train}})}^{\alpha}\left(Y^{\mathrm{test}}\mid\mathcal{C}(Y^{\mathrm{train}})\right)$ that satisfies
$\mathbb{P}\left(\theta_{\mathcal{S}(Y^{\mathrm{train}})}\not\in\mathrm{CI}_{\mathcal{S}(Y^{\mathrm{train}})}^{\alpha}\left(Y^{\mathrm{test}}\mid\mathcal{C}(Y^{\mathrm{train}})\right)\mid\mathcal{C}(Y^{\mathrm{train}})\right)\leq\alpha,$ (2.3)
@@ 2.1 The split-and-condition framework
for some conditioning set $\mathcal{C}(\cdot)$ such that $\sigma\left(\mathcal{C}(Y^{\mathrm{train}})\right)\supseteq\sigma\left(\mathcal{S}(Y^{\mathrm{train}})\right)$, where $\sigma\left(\cdot\right)$ denotes the sigma algebra generated by a random variable, i.e. the coverage guarantee in (2.3) conditions on more than that in (2.2). By the law of total probability, (2.3) implies (2.2), and both imply marginal coverage (Fithian et al., 2017).
Table 1 specifies the conditioning set $\mathcal{C}(\cdot)$ and the relationship between $Y^{\mathrm{train}}$, $Y^{\mathrm{test}}$, and $Y$ for some well-known split-and-condition approaches. Additional details are as follows:
[…]
    $\mathbb{P}\left(\theta_{\mathcal{S}(Y^{\mathrm{train}})}\not\in\mathrm{CI}_{\mathcal{S}(Y^{\mathrm{train}})}^{\alpha}\left(Y^{\mathrm{test}}\right)\right)\leq\alpha.$ (2.4)
  • In data fission (Leiner et al., 2023), once again $\mathcal{C}(Y^{\mathrm{train}})=Y^{\mathrm{train}}$, but now $Y^{\mathrm{train}}$ and $Y^{\mathrm{test}}$ are not constructed to be independent. Thus, (2.3) takes the form
    $\mathbb{P}\left(\theta_{\mathcal{S}(Y^{\mathrm{train}})}\not\in\mathrm{CI}_{\mathcal{S}(Y^{\mathrm{train}})}^{\alpha}\left(Y^{\mathrm{test}}\mid Y^{\mathrm{train}}\right)\mid Y^{\mathrm{train}}\right)\leq\alpha.$ (2.5)
  • In conditional selective inference (Fithian et al., 2017), we let $Y^{\mathrm{train}}=Y^{\mathrm{test}}=Y$. Thus, (2.3) takes the form
    $\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{C}(Y)}^{\alpha}\left(Y\mid\mathcal{C}(Y)\right)\mid\mathcal{C}(Y)\right)\leq\alpha,$ (2.6)
    for some $\sigma\left(\mathcal{C}(Y^{\mathrm{train}})\right)\supseteq\sigma\left(\mathcal{S}(Y^{\mathrm{train}})\right)$. Furthermore, the power of conditional selective inference can be greatly increased via randomization (Tian and Taylor, 2018; Panigrahi et al., 2024).
In general, constructing a confidence interval for data thinning and sample splitting (2.4) is no more complicated than constructing the classical confidence interval (2.1). Constructing a confidence interval for data fission (2.5) or conditional selective inference (2.6) is more challenging, and may require bespoke approaches tailored to the problem at hand.
Table 1: For four examples of the split-and-condition framework, details of $Y^{\mathrm{train}}$, $Y^{\mathrm{test}}$, and $\mathcal{C}(Y^{\mathrm{train}})$ in (2.3) are provided. Data thinning and data fission require introduction of a noise random variable $\zeta$. Additional details are provided in Section 2.1.
Method $Y^{\mathrm{train}}$ $Y^{\mathrm{test}}$ $\mathcal{C}(Y^{\mathrm{train}})$ Note
[…]
Data fission $g_{\text{train}}(Y,\zeta)$ $g_{\text{test}}(Y,\zeta)$ $Y^{\mathrm{train}}$
Conditional selective inference $Y$ $Y$ $S(Y)$
@@ 2.2 The infer-and-widen framework
We now consider the infer-and-widen framework, outlined in Box 1. This involves a modification to the classical confidence interval (2.1), to account for the fact that $\theta_{\mathcal{S}(Y)}$ is a data-driven parameter. Based on the particulars of the selection function $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$, a modified $\alpha$-level, $H(\alpha,S)$, is derived so that $\mathrm{CI}^{H(\alpha,S)}_{\mathcal{S}(Y)}(Y)$ is valid at level $\alpha$, i.e.
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{H(\alpha,S)}(Y)\right)\leq\alpha.$ (2.7)
@@
In a Bonferroni correction we have $H(\alpha,S)=\alpha/|\mathcal{I}|$, and in the Holm-Bonferroni correction it is slightly larger. Since these approaches do not adapt to the details of the selection rule $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$, the constant $H(\alpha,S)$ is easy to compute but overly conservative.
@@
The foundational PoSI proposal (Berk et al., 2013) considers the setting of linear regression with fixed regressors. They derive a constant $H(\alpha,S)$ that enables inference on the coefficient(s) in any submodel, regardless of the form of $\mathcal{S}(\cdot)$. However, for a particular selection rule $\mathcal{S}(\cdot)$, the resulting confidence interval may be quite conservative. Furthermore, computing the constant $H(\alpha,S)$ is challenging unless the number of regressors is quite small.
@@
We note that despite the notation $H(\alpha,S)$ used in (2.7), the aforementioned infer-and-widen proposals result in confidence intervals whose widths do not depend on the particulars of the selection rule $\mathcal{S}(\cdot)$: that is, $H(\alpha,S)$ could be written more simply as $H(\alpha)$.
By contrast, the recent algorithmic stability proposal of Zrnic and Jordan (2023) provides constants $H(\alpha,S)$ that are tailored to the particular selection rule $S(\cdot)$ under consideration, and therefore has the potential to yield narrow confidence intervals at a given nominal coverage. The notion of stability underlying algorithmic stability is as follows.
Definition 1 (Definitions 4.1 and 4.2, Zrnic and Jordan (2023)).
Two random variables $Q$ and $T$ are $(\eta,\tau)$-indistinguishable from each other, denoted $Q\approx_{(\eta,\tau)}T$, if for all measurable sets $\mathcal{O}$, it holds that
$\mathbb{P}(Q\in\mathcal{O})\leq e^{\eta}\mathbb{P}(T\in\mathcal{O})+\tau.$ (2.8)
A (possibly randomized) selection rule $S:\mathcal{Y}\mapsto\mathcal{I}$ is $(\eta,\tau,\nu)$-stable with respect to a distribution $\mathcal{P}$ on $\mathcal{Y}$ if there exists a (possibly random, possibly dependent on $\mathcal{P}$) variable ${S}_{0}$ such that
$\mathbb{P}\left(\mathcal{S}(Y)\approx_{(\eta,\tau)}{S}_{0}\right)\geq 1-\nu.$ (2.9)
Since $S_{0}$ is independent of $Y$, we can use the classical interval (2.1) to conduct inference on a target given by $S_{0}$, i.e. it holds that $\mathbb{P}\left(\theta_{S_{0}}\not\in\mathrm{CI}_{S_{0}}^{\alpha}(Y)\right)\leq\alpha$. The key insight underlying algorithmic stability is that if $\mathcal{S}(Y)$ is stable, then for most $Y\sim\mathcal{P}$, the distribution of $\mathcal{S}(Y)$ is quite close to that of $S_{0}$: and so it is “almost as if” the selection rule $\mathcal{S}(Y)$ did not involve $Y$. This notion is formalized in Theorem 2.
Theorem 2 (Theorem 5.2, Zrnic and Jordan (2023)).
Let $\mathcal{S}(\cdot)$ be a $(\eta,\tau,\nu)$-stable selection rule with parameters possibly dependent on the data distribution $\mathcal{P}$ of $Y$. Assume that for any fixed $\theta_{i}\in\Theta$, we have access to a confidence interval $\mathrm{CI}_{i}^{\alpha}(Y)$ for $\theta_{i}$ with coverage at least $1-\alpha$, i.e. it satisfies (2.1). Then
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{(\alpha-\tau-\nu)e^{-\eta}}(Y)\right)\leq\alpha.$ (2.10)
Thus, to conduct inference on a data-driven target $\theta_{\mathcal{S}(Y)}$, we can apply the infer-and-widen interval in (2.7) with $H(\alpha,S)=(\alpha-\tau-\nu)e^{-\eta}$, a function of the stability of the selection rule $\mathcal{S}(\cdot)$. The values of $\eta,\tau,\nu$ that govern the stability of $\mathcal{S}(\cdot)$ directly determine the width of the confidence interval: the smaller these parameters, the narrower the confidence interval. Zrnic and Jordan (2023) construct stable selection rules in a number of settings via the introduction of auxiliary randomness. For technical reasons, they make use of Laplacian randomness; however, in the following sections we will explore both Gaussian and Laplacian auxiliary randomness to facilitate our investigation of infer-and-widen versus split-and-condition.
@@
We close with a remark about the infer-and-widen framework: at first glance, it appears that an “advantage” of infer-and-widen, over (for instance) conditional selective inference, is that the interval in (2.7) is of the same form as the classical interval in the absence of selection (2.1); i.e., no bespoke methods for computing a confidence interval are required. However, this is a double-edged sword, in the sense that bias induced by the selection event is perpetuated into the confidence interval. For instance, in the linear regression setting considered by Berk et al. (2013), the confidence intervals are centered around the coefficient estimates for the selected model, and hence their midpoints are highly biased. Thus, an infer-and-widen confidence interval that attains the nominal coverage must necessarily be extremely wide, to account for bias in the interval’s midpoint. We will see in the coming sections that this is a major drawback of infer-and-widen, and leads to intervals that are substantially wider than those arising from much simpler split-and-condition alternatives.

### v2

# Infer-and-widen, or not?

@@ Abstract
In recent years, there has been substantial interest in the task of selective inference: inference on a parameter that is selected from the data. Many of the existing proposals fall into what we refer to as the infer-and-widen framework: they produce symmetric confidence intervals whose midpoints do not account for selection and therefore are biased; thus, the intervals must be wide enough to account for this bias. In this paper, we investigate infer-and-widen approaches in three vignettes: the winner’s curse, maximal contrasts, and inference after the lasso. In each of these examples, we show that a state-of-the-art infer-and-widen proposal leads to confidence intervals that are much wider than simple alternatives when all methods are tuned to yield identical randomized selection events. Furthermore, even an “oracle” infer-and-widen confidence interval — the narrowest possible interval that could be theoretically attained via infer-and-widen — is often wider than these alternatives.
@@ 1 Introduction
In classical statistics, the target of inference — for instance, the null hypothesis that we wish to test, or the parameter for which we wish to construct a confidence interval — is fixed. However, motivated by the explosion of data across a variety of fields, recent interest has focused on the setting where the target of inference is itself a function of the data, and therefore is random. The terms selective inference and post-selection inference describe the set of strategies available for conducting inference on a random target.
@@
To formalize this, consider a random variable $Y$ with domain $\mathcal{Y}$, and a collection of parameters $\theta$ whose elements can be indexed by a (possibly infinite) index set $\mathcal{I}$. For a fixed index $i\in\mathcal{I}$, and for $\alpha\in(0,1)$, a classical confidence interval for $\theta_{i}$ is a confidence interval $\mathrm{CI}_{i}^{\alpha}(Y)$ with $1-\alpha$ coverage in the sense that
@@
Now, suppose that $i\in\mathcal{I}$ is no longer a fixed index; instead, some selection rule ${\mathcal{S}}:\mathcal{Y}\to\mathcal{I}$ uses the data to select a parameter on which to conduct inference. A confidence interval provides valid unconditional selective coverage (Fithian et al., 2017) at level $\alpha$ if
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{\alpha}(Y)\right)\leq\alpha.$ (1.2)
Unconditional coverage is not guaranteed by classical confidence intervals. In this paper, we consider methods that yield unconditional guarantees, in the sense of 1.2. Some of these confidence intervals also yield (stronger) conditional selective guarantees, in the sense that
$\mathbb{P}\left(\theta_{i}\not\in\mathrm{CI}_{i}^{\alpha}(Y)~{}|~{}\mathcal{C}(Y)\right)\leq\alpha$ (1.3)
@@
for an event $\mathcal{C}(Y)$ satisfying $\sigma\left(\mathcal{C}(Y)\right)\supseteq\sigma\left({\mathcal{S}}(Y)\right)$, where $\sigma(\cdot)$ denotes the sigma algebra generated by a random variable, i.e., the coverage guarantee in 1.3 conditions on at least the selection event; by the law of total expectation, conditional coverage 1.3 implies unconditional coverage 1.2.
@@ 1.1 The infer-and-widen framework
As presented in Box 1.1, the infer-and-widen framework involves (Step 1) defining a target of inference using some or all of the data, and then (Step 2) symmetrically adjusting the critical values arising from “classical” inference to account for the fact that the target of inference is not fixed in order to achieve unconditional coverage (1.2). The simplest examples of this strategy are the Bonferroni (Bonferroni, 1936), Holm-Bonferroni (Holm, 1979), and Šidák (Šidák, 1967) procedures, which are suitable when a finite number of targets of inference are considered (and in the case of Šidák, under additional assumptions). In the special case of inference on contrasts of a linear model, the Scheffé correction is available (Scheffé, 1959); this idea is further refined by the PoSI method (Berk et al., 2013). While conceptually attractive, these approaches tend to be quite conservative (Bonferroni, Holm-Bonferroni, Šidák), are limited to a linear model setting (Scheffé, PoSI), or are computationally intractable (PoSI).
Other recent infer-and-widen approaches have attempted to address these shortcomings. The local simultaneous inference approach of Zrnic and Fithian (2024b) improves upon the Bonferroni correction by focusing on the subset of hypotheses that are most likely given the observed data. By contrast, the algorithmic stability proposal of Zrnic and Jordan (2023) adapts to the details of the selection algorithm. By drawing connections with the field of differential privacy (Dwork and Roth, 2014), the “stability” of the approach for selecting the target of inference in Step 1 of Box 1.1 is quantified; this, in turn, leads to a stability-specific adjustment of the critical values in Step 2 of Box 1.1. In a sense, algorithmic stability achieves the lofty goals laid out in the PoSI proposal of Berk et al. (2013) by creating a framework to compute a selection-specific adjustment to the critical values (see also the online supplement).
[!htb] Box 1: Infer-and-widen. 1. A target of inference is selected from the data. 2. “Classical” inference is conducted on all of the data, with a symmetric adjustment of the critical values to account for the data-driven selection of the target of inference.
1.2 Other approaches for valid selective inference
Not all approaches that yield unconditional guarantees of the form (1.2) fall into the infer-and-widen framework. Some methods produce asymmetric confidence intervals around the biased point estimate: for instance, Benjamini et al. (2019) and Fuentes et al. (2018) construct an asymmetric Bonferroni correction that accounts for the direction of the bias. More recently, Zrnic and Fithian (2024a) extend the local simultaneous approach of Zrnic and Fithian (2024b) by additionally adapting the correction to the specific selection algorithm. The hybrid approaches of Andrews et al. (2019) and McCloskey (2024) apportion a fraction of the Type 1 error budget to rule out unlikely selection events, and then construct a conditionally-valid confidence interval on the likely hypotheses; note that this only guarantees unconditional coverage 1.2.
Furthermore, a number of methods yield coverage guarantees conditional on the selection event, i.e., (1.3). Some methods do this by obtaining independent training and test sets; then, the training set can be used to define the target of inference, and the test set to conduct inference on this target. Because the two sets are independent, we can conduct inference using the test set as though the target of inference were fixed, despite the fact that it is actually a function of the training data. Sample splitting, applicable in the setting of independent and identically distributed observations, provides one approach to obtain independent training and test sets (Cox, 1975). More recent data thinning proposals generalize the idea of sample splitting, and can be applied when only a single observation is available (Rasines and Young, 2023; Neufeld et al., 2024; Dharamshi et al., 2024).
In cases where it is either impossible or undesirable to create independent training and test sets, strategies that achieve conditional guarantees of the form 1.3 may still be available. For instance, a target of inference can be defined using all of the data, and then inference can be conducted conditional on the event that the data led to this particular target of inference. Examples include the proposals of Fithian et al. (2017); Lee et al. (2016); Chen and Witten (2023); Gao et al. (2024); Andrews et al. (2019). Alternatively, if a researcher is willing to define the target of inference using only a subset of the data or using a noisy version of the data, then data fission Leiner et al. (2023) or randomized conditional selective inference Tian and Taylor (2018); Panigrahi et al. (2024) may be applicable.
1.3 Contributions and organization
In this paper, we conduct an empirical and theoretical evaluation of infer-and-widen strategies. Since randomization is known to improve the performance of selective inference methods (Panigrahi et al., 2024; Tian and Taylor, 2018), we only consider randomized infer-and-widen approaches. The bottom line of our paper is as follows: when comparing infer-and-widen strategies to alternative selective inference methods that make use of identical selection events and yield conditional 1.3 (rather than unconditional 1.2 guarantees), the latter often outperform the former in terms of confidence interval width. Furthermore, the latter often outperform even the “oracle” infer-and-widen procedure, the narrowest possible interval that could be theoretically attained within the infer-and-widen framework.
The rest of this paper is organized as follows. In Section 2, we formalize the infer-and-widen framework. In Sections 3, 4, and 5, we compare its performance to alternative approaches in two vignettes considered in Zrnic and Jordan (2023), and in an example involving model selection via the lasso. In Section 6, we show analytically that the infer-and-widen interval can suffer from a highly biased midpoint. We close with a discussion in Section 7. Theoretical results are proven in the online supplement. Scripts to fully reproduce all numerical results can be found at https://github.com/rflperry/infer_and_widen.
2 Formalizing the infer-and-widen framework
@@
We now formalize the infer-and-widen (I&W) framework, outlined in Box 1.1. This involves a modification to the classical confidence interval $\mathrm{CI}_{i}^{\alpha}$ in (1.1) to account for the fact that $\theta_{\mathcal{S}(Y)}$ is a data-driven parameter. Based on the particulars of the selection function $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ and data $Y$, a modified $\alpha$-level, $H(\alpha,{\mathcal{S}},Y)$, is derived so that $\mathrm{CI}^{H(\alpha,{\mathcal{S}},Y)}_{\mathcal{S}(Y)}(Y)$ is unconditionally valid at level $\alpha$ in the sense of 1.2, i.e.,
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{H(\alpha,{\mathcal{S}},Y)}(Y)\right)\leq\alpha.$ (2.1)
In effect, this procedure symmetrically widens the classical confidence interval to account for bias in its midpoint.
@@
In a Bonferroni correction we have $H(\alpha,{\mathcal{S}},Y)=\alpha/|\mathcal{I}|$, and in the Holm-Bonferroni correction it is slightly larger. Since these approaches adapt neither to the selection rule $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ nor to the data $Y$, the constant $H(\alpha,{\mathcal{S}},Y)$ is easy to compute but overly conservative.
@@
The foundational PoSI proposal (Berk et al., 2013) considers the setting of linear regression with fixed regressors. They derive a constant $H(\alpha,{\mathcal{S}},Y)$ that enables inference on the coefficient(s) in any submodel, regardless of the form of $\mathcal{S}(\cdot)$ or observed data $Y$. However, for a particular selection rule $\mathcal{S}(\cdot)$, the resulting confidence interval may be quite conservative, and computing the constant $H(\alpha,{\mathcal{S}},Y)$ is challenging or infeasible.
@@
Despite the notation $H(\alpha,{\mathcal{S}},Y)$ used in (2.1), the aforementioned infer-and-widen proposals result in confidence intervals whose widths do not depend on the particulars of the selection rule $\mathcal{S}(\cdot)$ nor data $Y$: that is, $H(\alpha,{\mathcal{S}},Y)$ could be written more simply as $H(\alpha)$.
In contrast are two newer approaches. The local simultaneous inference method of Zrnic and Fithian (2024b) apportions the Type-1 error budget $\alpha$, first to use the data $Y$ to rule out the most unlikely hypotheses, and then to perform a Bonferroni correction over a subset of all $\left|\mathcal{I}\right|$ hypotheses. We note that in this case, $H(\alpha,{\mathcal{S}},Y)$ could be written more simply as $H(\alpha,Y)$, as the correction does not take into account the selection rule. On the other hand, the algorithmic stability proposal of Zrnic and Jordan (2023) provides constants $H(\alpha,{\mathcal{S}})$ that are tailored to the particular selection rule ${\mathcal{S}}(\cdot)$ under consideration, although not to the data $Y$. The online supplement contains an overview of the algorithmic stability proposal; intuitively, a more “stable” selection rule ${\mathcal{S}}(\cdot)$ requires a less substantial adjustment to the critical values in Step 2 of Box 1.1.
@@
We close with a remark about the infer-and-widen framework: at first glance, it appears that an “advantage” of infer-and-widen, over alternative selective inference approaches, is that the interval in (2.1) is of the same form as the classical interval in the absence of selection (1.1), i.e., no bespoke methods for computing a confidence interval are required. However, this is a double-edged sword, in the sense that bias induced by the selection event is perpetuated into the confidence interval. Thus, an infer-and-widen confidence interval that attains the nominal coverage may be extremely wide, to account for bias in the interval’s midpoint. We will see that this is a major drawback of infer-and-widen.